# Gaussian Building Mesh (GBM): Extract a Building’s 3D Mesh with Google Earth and Gaussian Splatting

### v2

@@ V-C View Synthesis
V-E Discussions
@@
VI Conclusion
By utilizing Google Earth Studio and integrating SAM2/GroundingDINO, 2DGS+, and our mask refinement algorithm, we developed a robust pipeline capable of extracting a colored 3D mesh of a building using its address, postal code, geographic coordinates, or location name. We benchmarked our method against GS2Mesh, which, to the best of our knowledge, is the only comparable segmentation-based Gaussian Splatting meshing algorithm publically available. Our results demonstrate that our approach produces significantly higher quality meshes and is significantly less prone to failure.
Acknowledgments
We would like to thank the authors of 2DGS [27] and GS2Mesh [2], and the Github contributors of 2DGS+ [5] for making their codes publicly available.

### v3

@@ 5.3 View Synthesis
5.5 Ablation Studies
Table 4: Ablation Study on 2DGS+ for View Synthesis and 3D Meshing
[…]
The ablation study is performed on a separate Westminster Palace scene comparing the improvements of 2DGS+ to 2DGS as a baseline. A mip-NeRF (Barron et al., 2021) style training/testing split was used. Both view synthesis and 3D mesh quality assessments are used. Additionally, training time was measured. Trim-Gaussian’s (Fan et al., 2024) and PixelGS’s (Zhang et al., 2024) contribution-based Gaussian pruning is denoted as Trim. Progressive propagation from GaussianPro(Cheng et al., 2024) is denoted as Prog. The standard 2DGS normal-based supervision is denoted as Norm. The Edge-Aware Normal/Geometry Supervision from AtomGS (Liu et al., 2024) is denoted EANorm. The Appearance Network from GOF (Yu et al., 2024) is denoted as Appearance. Fast-SSIM from Taming-GS (Mallick et al., 2024) is denoted as Fast. Doubled training iteration is denoted as 60k.
5.6 Discussions
@@
6 Conclusion
The ability to generate high-quality 3D building models from minimal user input addresses a key need in digital building monitoring workflows. We developed a robust pipeline capable of automatically extracting a colored 3D mesh of a building using its address, postal code, geographic coordinates, or location name directly. Our pipeline does not require image, video, or point cloud data from the user, and instead extracts the data from Google Earth Studio. We benchmarked our method against GS2Mesh, which, to the best of our knowledge, is the only comparable segmentation-based Gaussian Splatting meshing algorithm publicly available. Our results demonstrate that our approach produces significantly higher-quality meshes and is significantly less prone to failure, allowing for automated extraction of the 3D mesh of any user-designated buildings given Google Earth image coverage. These capabilities open the door to scalable applications in digital twin creation, construction monitoring and automated inspection, disaster monitoring, and many other downstream tasks.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
This work was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant (No. RGPIN-2022-03741).
We would like to thank the authors of 2DGS (Huang et al., 2024) and GS2Mesh (Wolf et al., 2025), and the GitHub contributors of 2DGS+ (Ye and GitHub-Contributors, 2024) for making their codes publicly available.